### v1

@@ 3.3 Evaluation Metrics
where $div$ is the distance of the counterfactuals $x^{*}$ from the factuals $x$, $S_{x}$, $S_{x^{*}}$ are the sets of all distances of the factual image to the images that have the same label as the factual and the counterfactual, respectively: $S_{x}=\{d(x,x^{\prime})|pa_{x^{\prime}}=pa_{x}\}$, $S_{x^{*}}=\{d(x,x^{\prime})|pa_{x^{\prime}}=pa_{x^{*}}\}$, where $x^{\prime}$ is any image from the data distribution and $pa_{x^{\prime}}$ is the variable being intervened upon. To compute the distance $d(\cdot)$, we train an unconditional VAE and use the KL divergence between the latent distributions: $d(x_{i},x_{j})=D_{KL}(\mathcal{N}(\mu_{i},\sigma_{i}),\mathcal{N}(\mu_{j},\sigma_{j})$. This metric is minimised by keeping both probabilities low, which represent a trade-off between being far from the factual class but not as far as other (real) images from the counterfactual class are. The weights $w_{1}$, $w_{2}$ depend on the effective difference each variable has on the image.
(a) MorphoMNIST
t: thickness, i: intensity, d: digit, im: image
(b) CelebA
s: smiling, e: eyeglasses, im: image
Figure 1: Causal graphs for our datasets, MorphoMNIST and CelebA.
4 Results
Datasets Given the availability of results and architectures from literature, we benchmark all the above models on MorphoMNIST (32x32) [1] and CelebA (64x64) [26]. We assume the causal graphs of Figure 1 (a) for MorphoMNIST and (b) for CelebA. MorphoMNIST is a purely synthetic dataset generated by inducing morphological operations on the well-established MNIST digits. We note, that interventions on thickness affect both the intensity and image (Figure 1(a)). Therefore, we should be able to measure this effect on intensity by comparing the prediction of the normalising flow for this variable with the prediction of the regressor for the produced image. Meanwhile, CelebA is based on real world images, and hence we use two assumed disentangled attributes as done in [30] Figure 1(b). No additional layer of parents is involved. These attributes are independent to each other which satisfies the criterion for counterfactual inference.
Setup For Normalising Flows we leverage the normflows package [44], using the QuadraticSpline, ConstScaleShift and conditional ScaleShift flow layers of [22]. For the MorphoMNIST dataset, we compare the VAE architecture as given in the original Deep-SCM paper [31], the HVAE architecture of [4] and the fine-tuned GAN architecture of [45] as the set up of [3] failed to converge. For CelebA, we use the VAE architecture proposed as baseline in [3], while we extend the HVAE used in [30] with the conditioning mechanisms of [4]. Finally, for the GAN we used the architecture proposed in [3].
We compare VAE with a fixed standard deviation of $\sqrt{e^{-5}}$ as used in [31, 3] versus a VAE with learnable standard deviation. For MorphoMNIST we trained the HVAE as described in 3.2 and used the counterfactual training for CelebA as we found it to ignore conditioning, detailed in 4.2. We experimented with different values for $\beta$ and found that $\beta=1$ and $\beta=5$ perform better for MorphoMNIST and CelebA, respectively. For GAN, we found that fine-tuning the cyclic cost minimisation was necessary to ensure closeness to the factual image for both datasets.
Table 1: Composition for MorphoMNIST. For embeddings, we concatenate the last layer activations across all predictors trained.
$l_{1}$ image space $\downarrow$ $l_{1}$ embeddings $\downarrow$ LPIPS $\downarrow$
Model 1 cycle 10 cycles 1 cycle 10 cycles 1 cycle 10 cycles
VAE ($\sigma=\sqrt{e^{-5}}$) $4.244_{2.158}$ $9.855_{4.768}$ $0.095_{0.042}$ $0.158_{0.064}$ $0.039_{0.003}$ $0.098_{0.003}$
VAE ($\sigma$ learned) $2.600_{1.454}$ $7.698_{4.199}$ $0.064_{0.033}$ $0.135_{0.057}$ $0.026_{0.003}$ $0.075_{0.003}$
HVAE $\mathbf{0.438_{0.178}}$ $\mathbf{1.550_{0.541}}$ $\mathbf{0.009_{0.004}}$ $\mathbf{0.038_{0.015}}$ $\mathbf{0.006_{0.001}}$ $\mathbf{0.024_{0.003}}$
GAN $16.779_{8.406}$ $16.780_{8.401}$ $0.193_{0.065}$ $0.193_{0.065}$ $0.154_{0.004}$ $0.154_{0.004}$
GAN fine-tuned $3.807_{1.822}$ $11.697_{5.750}$ $0.089_{0.042}$ $0.272_{0.106}$ $0.049_{0.004}$ $0.184_{0.008}$
4.1 Composition
To quantitatively evaluate composition, we perform a null intervention. Following the protocol of [30], we apply composition for one and for ten cycles, measuring the distance between the initial observable image and its first and tenth reconstructed versions respectively. In addition to the $\ell_{1}$ distance in the pixel space, we also use the LPIPS [53] metric and the $\ell_{1}$ in embedding space.
(a) VAE
(b) HVAE
(c) GAN fine-tuned
Figure 2: Composition for 10 cycles on the MorphoMNIST test set.
MorphoMNIST: In Table 1 we show quantitative results for composition. We extract embeddings by concatenating the last layers of the anticausal predictors and calculate LPIPS on VGG. We can observe that HVAE significantly outperforms all models across metrics both for $composition^{1}$ and $composition^{10}$, while GAN without cyclic cost minimisation performs poorly even from the first cycle. Figure 2 depicts qualitative results for the 10-cycle composition. It is evident that the HVAE is capable of preserving extremely well the details of the original image; the VAE performs adequately; whereas the GAN introduces the largest distortion after multiple cycles.
Table 2: Composition for CelebA. For the embeddings we use the last layer activation of a pretrained VGG-16 network.
$l_{1}$ image space $\downarrow$ $l_{1}$ embeddings $\downarrow$ LPIPS $\downarrow$
Model 1 cycle 10 cycles 1 cycle 10 cycles 1 cycle 10 cycles
VAE ($\sigma=\sqrt{e^{-5}}$) $126.8_{19.3}$ $126.4_{22.1}$ $0.536_{0.100}$ $0.553_{0.099}$ $0.332_{0.004}$ $0.462_{0.004}$
VAE ($\sigma$ learned) $127.8_{18.1}$ $\mathbf{123.4_{22.1}}$ $0.527_{0.102}$ $0.549_{0.102}$ $0.295_{0.004}$ $0.424_{0.005}$
HVAE fine-tuned $129.4_{11.6}$ $138.8_{23.4}$ $\mathbf{0.498_{0.121}}$ $\mathbf{0.525_{0.121}}$ $\mathbf{0.063_{0.003}}$ $\mathbf{0.200_{0.008}}$
GAN $118.9_{21.1}$ $126.1_{17.1}$ $0.548_{0.121}$ $0.660_{0.132}$ $0.403_{0.004}$ $0.506_{0.004}$
GAN fine-tuned $\mathbf{115.3_{22.0}}$ $128.4_{21.4}$ $0.516_{0.120}$ $0.592_{0.133}$ $0.290_{0.003}$ $0.462_{0.005}$
CelebA: We confirm empirically that computing distances on pixel space for a dataset that contains complex real world images is not informative and therefore we resort to pretrained VGG-16 embeddings and LPIPS to effectively capture meaningful differences. From Table 2 we observe that HVAE achieves the best composition scores in terms of $l^{(1)}$ embeddings and LPIPS distance, while GAN again performs the worst. We can further verify the efficacy of HVAE from the qualitative results of Figure 3. Particularly, it retains the details of the original image, while VAE progressively blurs and distorts it during repeating cycles. GAN with cyclic cost minimisation does not introduce blurriness, but alters significantly both the image style and content.
(a) VAE ($\sigma$ learned)
(b) HVAE
(c) GAN fine-tuned
Figure 3: Composition for 10 cycles on CelebA test set.
4.2 Effectiveness
MorphoMNIST: To quantitatively evaluate effectiveness on MorphoMNIST, we train convolutional regressors for continuous variables (thickness, intensity) and a classifier for categorical variables (digits). We perform random interventions on a single parent each time, as per [4]. In Table 3, we can observe that the mechanism of HVAE achieves the best performance in the majority of the interventional scenarios.
Table 3: Effectiveness on MorphoMNIST test set.
Model $do(t)$ $do(i)$ $do(y)$ $do(t)$ $do(i)$ $do(y)$ $do(t)$ $do(i)$ $do(y)$
VAE ($\sigma=\sqrt{e^{-5}}$) $0.125_{0.01}$ $0.346_{0.01}$ $0.138_{0.01}$ $9.382_{0.57}$ $20.642_{0.59}$ $11.279_{0.56}$ $0.988_{0.01}$ $\mathbf{0.989_{0.01}}$ $0.933_{0.02}$
VAE ($\sigma$ learned) $0.109_{0.01}$ $0.333_{0.02}$ $0.139_{0.01}$ $3.149_{0.26}$ $5.326_{0.29}$ $4.64_{0.35}$ $\mathbf{0.989_{0.01}}$ $0.988_{0.01}$ $0.775_{0.02}$
HVAE $\mathbf{0.086_{0.09}}$ $\mathbf{0.224_{0.03}}$ $\mathbf{0.117_{0.01}}$ $\mathbf{1.991_{0.18}}$ $\mathbf{3.517_{0.26}}$ $\mathbf{2.098_{0.17}}$ $0.985_{0.01}$ $0.935_{0.02}$ $\mathbf{0.972_{0.02}}$
GAN $0.227_{0.01}$ $0.631_{0.02}$ $0.214_{0.01}$ $6.680_{0.25}$ $9.795_{0.25}$ $6.632_{0.16}$ $0.810_{0.01}$ $0.811_{0.01}$ $0.813_{0.02}$
GAN fine-tuned $0.228_{0.01}$ $0.680_{0.02}$ $0.393_{0.02}$ $9.436_{0.59}$ $15.137_{0.99}$ $12.388_{0.57}$ $0.961_{0.02}$ $0.966_{0.01}$ $0.451_{0.024}$
We note that the error for intensity remains low when performing an intervention on thickness, compared to other interventions, which showcases the effectiveness of the conditional intensity mechanism. The compliance with the causal graph of Figure 1 can also be confirmed qualitatively in Figure 4. Intervening on thickness modifies intensity, while intervening on intensity does not affect thickness (graph mutilation). Interventions on the digit alter neither.
(a) VAE ($\sigma$ learned)
(b) HVAE
(c) GAN fine-tuned
Figure 4: From left to right: Factual, $do(thickness)$, $do(intensity)$, $do(digit)$
CelebA: To measure counterfactual effectiveness for CelebA, we train binary classifiers for the smiling and eyeglasses attributes. We observe that by training HVAE only with the standard ELBO loss, the model exclusively reconstructs the initial image and systematically ignores the causal/counterfactual condition. To alleviate this, we exploit the counterfactual training process that is proposed in [4]. In brief, the lowest ELBO loss is incorporated early during training into the counterfactual loss by introducing a Lagrange coefficient, thus transforming the setup into a Lagrangian optimisation problem [33].
Table 4 presents the measured effectiveness in terms of F1 score using the attribute classifiers. Reasonable F1 values indicate that all models [^4] HVAE without counterfactual training was unaffected by interventions, so it was not included in Table 4 are capable of performing counterfactual predictions. The fine-tuned HVAE outperforms all other models across most scenarios with the most evident difference being observed in interventions on the attribute eyeglasses. This suggests that HVAE can efficiently manipulate this attribute and generate plausible counterfactuals. On the other hand, we notice that when we intervene on the attribute smiling, the fine-tuned GAN outperforms other models, which indicates that, even if it cannot perform accurate interventions, it can better disentangle this intervention from the attribute eyeglasses.
The following standard data augmentations were applied to mitigate biases existing in the CelebA dataset, which were exploited by the classifiers during training: random flipping, random cropping and colour distortions. Additionally, as CelebA is a highly unbalanced dataset, a weighted sampler is used to ensure a balanced class distribution in each training batch. Arguably, effectiveness is challenging to quantify, since the data-dependent methods that are typically used may lead to biased measurements.
Table 4: Effectiveness on CelebA test set.
Smiling (s) F1. $\uparrow$ Eyeglasses (e) F1. $\uparrow$
Model $do(s)$ $do(e)$ $do(s)$ $do(e)$
VAE ($\sigma=\sqrt{e^{-5}}$) $0.981_{0.01}$ $0.996_{0.01}$ $0.929_{0.05}$ $0.771_{0.02}$
VAE ($\sigma$ learned) $0.897_{0.02}$ $0.987_{0.01}$ $0.938_{0.05}$ $0.810_{0.02}$
HVAE fine-tuned $\mathbf{0.998_{0.01}}$ $\mathbf{0.997_{0.01}}$ $0.883_{0.06}$ $\mathbf{0.981_{0.02}}$
GAN $0.818_{0.02}$ $0.837_{0.02}$ $0.824_{0.04}$ $0.873_{0.01}$
GAN fine-tuned $0.819_{0.02}$ $0.873_{0.01}$ $\mathbf{0.957_{0.03}}$ $0.891_{0.01}$
In Figure 5 we observe the apparent weakness of VAEs to perform on more complex, real-world images, since its single latent variable can capture only few details from the original images, introducing bluriness. Due to the expressivity of the hierarchical latents, the fine-tuned HVAE is able to retain much of the information of the original image and effectively modify it by following the counterfactual condition. The fine-tuned GAN is also capable of producing counterfactuals that are similar to the factual and adhere to the condition, but of lower quality compared to the HVAE.
(a) VAE ($\sigma$ learned)
(b) HVAE fine-tuned
(c) GAN fine-tuned
Figure 5: From left to right: Factual, $do(Smiling)$, $do(Eyeglasses)$
4.3 Realism & Minimality
The FID metric in Table 5 aligns with our qualitative results, showing that HVAE outperforms VAEs on MorphoMNIST and GANs on CelebA. HVAE without counterfactual training performs better on CelebA, as expected, since it does not introduce any information.
For minimality, we report the weighted CLD metric and its component probabilities (Table 5). We report lower values for VAEs on MorphoMNIST [^5] We discretised continuous variables using 10 bins., as shown both in CLD and $P(S_{x^{*}}\leq div)$, with HVAE following. In CelebA, HVAEs outperform both VAEs and fine-tuned GAN, whose scores are comparable. We notice that for CelebA, the term $P(S_{x}\geq div)$ cannot effectively discern changes that are too minimal from effective ones, which can be explained by the fact that the intervened variables are not as distinctive as others such as gender or background, making the factual-counterfactual distance essentially too small. This is the motivation for the weighting terms on CLD, which for these experiments are $w_{1}=0.8$, $w_{2}=0.2$. We trained unconditional VAEs for both datasets and found the metric performing similarly even when using distance between different types of embeddings, such as VGG-16 or CLIP [34].
Table 5: Realism and Minimality for MorphoMNIST and CelebA. In the minimality columns, the numbers from left to right are the values for: (i) weighted CLD, (ii) $P(S_{x^{*}}\leq div)$, (iii) $P(S_{x}\geq div)$.
MorphoMNIST CelebA
Model FID $\downarrow$ Minimality $\downarrow$ FID $\downarrow$ Minimality $\downarrow$
VAE ($\sigma=\sqrt{e^{-5}}$) $17.096$ $0.276$ $0.065$ $0.843$ $77.307$ $0.310$ $0.072$ $0.925$
VAE ($\sigma$ learned) $10.124$ $\mathbf{0.268}$ $0.055$ $0.839$ $66.412$ $0.297$ $0.014$ $0.984$
HVAE $\mathbf{5.362}$ $0.272$ $0.112$ $0.736$ $\mathbf{12.667}$ $0.295$ $0.0001$ $0.999$
HVAE fine-tuned - - $22.047$ $\mathbf{0.295}$ $0.003$ $0.996$
GAN $72.766$ $0.390$ $0.339$ $0.570$ $41.536$ $0.362$ $0.224$ $0.777$
GAN fine-tuned $35.568$ $0.286$ $0.119$ $0.764$ $31.560$ $0.304$ $0.046$ $0.953$
5 Discussion and Conclusions
Many methods have been introduced for counterfactual image generation. However, to date, a systematic way to assess current and potentially upcoming methods under a single comprehensive framework is missing. Given the rapidly emerging interest in causal representation learning and counterfactual generation, it is of paramount importance to build a unified framework to evaluate and cross-compare different approaches. Our aspiration is therefore to help as well as to inspire further work in the community on this direction. We evaluated three different setups for image mechanisms on four different evaluation metrics. We identify that the development of Deep-SCM-based conditional HVAE is in the right direction, as it outperformed all other models examined (VAEs, GANs) across all metrics performed. Our findings indicate that the hierarchical structure of the latents allows for more expressivity and accurate abduction of the noise variable, versus VAEs and GANs. Moreover, we created a user-friendly framework for follow-up work by future modellers, adaptable to a diverse set of causal methods and assessment metrics for the community to build on and expand further.
Limitations and future work: We examined three families of generative models and two datasets, under the Deep-SCM paradigm. Therefore, we purposely left methods that differ from this paradigm for future work, such as deep twin networks and backtracking counterfactuals. As future work, we also believe that it is important to investigate diffusion models (DM) which have showcased impressive performance on image generation. DM have not been conditioned yet with SCM graphs and we seek to explore this bold direction. An important consideration for DM is that abduction of noise is not always straightforward and thus, conditioning must methodically be considered. We also aim to explore the effect of more complex Deep-SCM structures on the models presented. Applying the proposed framework to further models, datasets and causal mechanisms will shed light on the distinct abilities of conditional generative models. It will also help the community to better understand how generative models interact with causal mechanisms to disentangle confounding attributes and conform to causal paths that determine image generation.
Acknowledgments
This work was partly supported by the National Recovery and Resilience Plan Greece 2.0, funded by the European Union under the NextGenerationEU Program (Grant: MIS 5154714). S.A. Tsaftaris acknowledges support from the Royal Academy of Engineering and the Research Chairs and Senior Research Fellowships scheme (grant RCSRF1819\8\25), and the UK’s Engineering and Physical Sciences Research Council (EPSRC) support via grant EP/X017680/1, and the UKRI AI programme, and the EPSRC, for CHAI - EPSRC AI Hub for Causality in Healthcare AI with Real Data [grant number EP/Y028856/1]. P. Sanchez thanks additional financial support from the School of Engineering, the University of Edinburgh.
Appendix
In Appendix 0.A we provide more details regarding the model architectures and hyperparameters for all the experiments we ran. In Appendix 0.B we include extra qualitative results for all the models we compare.
Appendix 0.A Experimental details
In the following tables that depict the architectures for our models, F refers to number of filters, K refers to kernel width and height, S refers to stride, P refers to padding, Up refers to the scale of upsampling, BN refers to batch normalization, D refers to dropout probability and A refers to activation function.
0.A.1 Normalising Flows
For Normalising Flows we leverage the normflows package [44] and base our implementation in [22]. While flows can be trained for all attribute variables, in practice we only employ them for the intensity variable, as we discussed in the main paper. In Table 6 we present the architecture for this flow, which is conditioned on its parent variable, thickness. For training the model we use the Adam optimizer with learning rate $10^{-3}$, batch size of 64 and early stopping with the patience parameter set to 2.
Table 6: Normalising Flow Architecture for intensity (MorphoMNIST).
Layers
Masked Affine Autoregressive Flow
Autoregressive Rational Quadratic Spline Flow
Autoregressive Rational Quadratic Spline Flow
Autoregressive Rational Quadratic Spline Flow
Sigmoid Transformation Flow
Constant AddScale Flow
Affine Coupling Flow
0.A.2 VAE
We adopt the VAE implementation as defined in [31, 4]. The output of the VAE decoder is passed though a Gaussian network that is implemented as a convolutional layer to obtain the corresponding $\mu$, $\sigma$. Then, the reconstructed image is sampled with the reparameterisation trick. To condition the model we concatenate the parent variables with the output features of the first Linear layer of the encoder (Tables 7 and 9).
0.A.2.1 MorphoMNIST
The architectures of the VAE encoder and decoder are displayed in Tables 7 and 8, respectively. We train the model until convergence (around 1M iterations) with batch size 256 using the AdamW optimizer with initial learning rate $10^{-3}$, $\beta_{1}$=0.9, $\beta_{2}$=0.9 and weight decay 0.01. Furthermore, we use early stopping with the patience parameter set to 10 and a linear warmup scheduler for the first 100 iterations. The dimension of the latent variable is set to 16. We also apply gradient clipping with value 350 and gradient skipping with value 500.
Table 7: VAE Encoder Architecture for MorphoMNIST.
Layer F K S P BN A
Conv2D 32 5 2 1 N ReLU
Conv2D 32 3 2 1 N ReLU
Conv2D 32 3 2 1 N ReLU
Linear 128 - - - N ReLU
Linear 128 - - - N ReLU
Linear (latent dim) 16 - - - N ReLU
Table 8: VAE Decoder Architecture for MorphoMNIST.
Layer F K S Up P BN A
Linear 128 - - - - N ReLU
Linear 512 - - - - N ReLU
Upsample - - - 2x - - -
Conv2D 32 3 1 - 1 N ReLU
Upsample - - - 2x - - -
Conv2D 32 3 1 - 1 N ReLU
Upsample - - - 2x - - -
Conv2D 16 5 1 - 2 N ReLU
0.A.2.2 CelebA
We extend the VAE implementation of [31, 4] to be compatible with higher resolution images (Tables 9 and 10). We train the model until convergence (around 600K iterations) with a batch size of 128 leveraging the Adam optimizer with initial learning rate $5\cdot 10^{-4}$ and $\beta_{1}$=0.9, $\beta_{2}$=0.999. We also use a linear warmup scheduler for the first 100 iterations. Additionally, we set the early stopping patience to 10 and add gradient clipping and gradient skipping with values 350 and 500 respectively. The dimension of the latent variable is set to 16.
Table 9: VAE Encoder Architecture for CelebA.
Layer F K S P D A
Conv2D 32 3 2 1 0.25 LReLU(0.01)
Conv2D 64 3 2 1 0.25 LReLU(0.01)
Conv2D 128 3 2 1 0.25 LReLU(0.01)
Conv2D 256 3 2 1 0.25 LReLU(0.01)
Conv2D 256 1 1 0 - LReLU(0.01)
Linear 256 - - - - LReLU(0.01)
Linear 256 - - - - LReLU(0.01)
Linear (latent dim) 16 - - - - LReLU(0.01)
Table 10: VAE Decoder Architecture for CelebA.
Linear F K S Up P BN D A
Linear 256 - - - - N - ReLU
Linear 4096 - - - - N - -
Upsample - - - 2x - - - -
Conv2D 256 3 1 - 1 Y 0.25 LReLU(0.01)
Upsample - - - 2x - - - -
Conv2D 128 3 1 - 1 Y 0.25 LReLU(0.01)
Upsample - - - 2x - - - -
Conv2D 64 3 1 - 1 Y 0.25 LReLU(0.01)
Upsample - - - 2x - - - -
Conv2D 32 3 1 - 1 Y 0.25 LReLU(0.01)
Upsample - - - 1x - - - -
Conv2D 16 3 1 - 1 Y - Sigmoid
0.A.3 HVAE
We follow the experimental setup of [4] leveraging the proposed conditional structure of the very deep VAE [2]. To condition the model we expand the parents and concatenate them with the corresponding latent variable at each top-down layer.
0.A.3.1 MorphoMNIST
: We use a group of 20 latent variables with spatial resolutions $\{1,4,8,16,32\}$. Each resolution includes 4 stochastic blocks. The channels of the feature maps per resolution are $\{256,128,64,32,16\}$, while the dimension of the latent variable channels is 16.
We train the model until convergence (approximately 1M iterations) with a batch size of 256 and apply early stopping with patience 10. The resolution of the input images is 32x32. We use the AdamW optimizer with initial learning rate $10^{-3}$, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.9, a weight decay of 0.01 and a linear warmup scheduler for the first 100 iterations. We additionally add gradient clipping and gradient skipping with values 350 and 500 respectively.
0.A.3.2 CelebA

### v5

@@ A.2 Metrics
where $div$ is the distance of the counterfactuals $x^{*}$ from the factuals $x$, $S_{x}$, $S_{x^{*}}$ are the sets of all distances of the factual image to the images that have the same label as the factual and the counterfactual, respectively: $S_{x}=\{d(x,x^{\prime})|pa_{x^{\prime}}=pa_{x}\}$, $S_{x^{*}}=\{d(x,x^{\prime})|pa_{x^{\prime}}=pa_{x^{*}}\}$, where $x^{\prime}$ is any image from the data distribution and $pa_{x^{\prime}}$ is the variable being intervened upon. To compute the distance $d(\cdot)$, we train an unconditional VAE and use the KL divergence between the latent distributions [^6] We computed CLD on the embedding space of a separately trained unconditional VAE for each dataset. We found other distances (LPIPS, CLIP) to perform similarly.: $d(x_{i},x_{j})=D_{KL}(\mathcal{N}(\mu_{i},\sigma_{i}),\mathcal{N}(\mu_{j},\sigma_{j})$. This metric is minimised by keeping both probabilities low, which represent a trade-off between being far from the factual class but not as far as other (real) images from the counterfactual class are. The weights $w_{1}$, $w_{2}$ depend on the effective difference each variable has on the image.
Appendix B Experimental Setup
B.1 Dataset preprocessing and splits
The train set of MorphMNIST consists of 48000 samples, while validation and test set consist of 12000 and 10000 samples, respectively. For the CelebA dataset, we used 162770 images for the training, while we kept 19867 images for the validation set and 19962 images for the counterfactual test set. For the ADNI dataset, after pre-processing (following the process of (59)) we get 2D axial grayscale slice images of 192x192 resolution. For each subject we use all visits (2-4 visits depending on the patient) and take the 20 middle MRI slices, ending up with a train set of 10780 images (180 subjects in total) and a test set of 2240 images (38 subjects).
B.2 Normalising Flows
For Normalising Flows we leverage the normflows package (51) and base our implementation in (27). We use the Adam optimizer with learning rate $10^{-3}$, batch size of 64 and early stopping with the patience parameter set to 10.
B.3 VAE
We adopt the VAE implementation as defined in (36; 7). The output of the VAE decoder is passed though a Gaussian network that is implemented as a convolutional layer to obtain the corresponding $\mu$, $\sigma$. Then, the reconstructed image is sampled with the reparameterisation trick. To condition the model we concatenate the parent variables with the output features of the first Linear layer of the encoder. We train the model until convergence with a batch size of 128 leveraging the Adam optimizer with initial learning rate $5\cdot 10^{-4}$ and $\beta_{1}$=0.9, $\beta_{2}$=0.999. We also use a linear warmup scheduler for the first 100 iterations. Additionally, we set the early stopping patience to 10 and add gradient clipping and gradient skipping with values 350 and 500 respectively. The dimension of the latent variable is set to 16.
B.4 HVAE
We follow the experimental setup of (7) leveraging the proposed conditional structure of the very deep VAE (5). To condition the model we expand the parents and concatenate them with the corresponding latent variable at each top-down layer.
For MorphoMNIST, we use a group of 20 latent variables with spatial resolutions $\{1,4,8,16,32\}$. Each resolution includes 4 stochastic blocks. The channels of the feature maps per resolution are $\{256,128,64,32,16\}$, while the dimension of the latent variable channels is 16.
We train the model until convergence (approximately 1M iterations) with a batch size of 256 and apply early stopping with patience 10. The resolution of the input images is 32x32. We use the AdamW optimizer with initial learning rate $10^{-3}$, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.9, a weight decay of 0.01 and a linear warmup scheduler for the first 100 iterations. We additionally add gradient clipping and gradient skipping with values 350 and 500, respectively.